ChrisTheFlash:

A woman who is homozygous dominant for Black Hair marries a man who is heterozygous for brown hair. What are the possible genotypes and phenotypes of their offspring be? * Your answer

QuestionCoveBot:

Do you know what homozygous and heterozygous mean?

ChrisTheFlash:

no

ChrisTheFlash:

this is just a question tho

QuestionCoveBot:

Homozygous: having two identical alleles of a particular gene or genes. Heterozygous: having two different alleles of a particular gene or genes.

ChrisTheFlash:

oh

ChrisTheFlash:

so what is the awnser

QuestionCoveBot:

The symbols black hair is: B The symbol for brown hair is: b

ChrisTheFlash:

ok i get it now but what is the answer

ChrisTheFlash:

hello

QuestionCoveBot:

For the Phenotypes: Make a punnet square B b _________ B | BB | Bb| B |BB | Bb| 1/2 BB (Black, Black) 1/2 Bb (Black, Brown) For the Genotypes: Make another punnet square B b _________ B | BB | Bb| B | BB | Bb| 1/2 BB (Black, Black) 1/2 Bb (Brown, Black)

ChrisTheFlash:

is this the answer

QuestionCoveBot:

No. I will get to that soon.

ChrisTheFlash:

then what is the answer

QuestionCoveBot:

So half the Phenotypes is BB. Half is Bb. And half the Genotypes is BB. Half is Bb. 1/2 1/2 1/2 1/2 Math time :) So, it's a higher chance for Black hair, since the husband is heterozygous. To start: Phenotype: BB 1/2 the time Bb 1/2 the time Genotype: BB 1/2 the time Bb 1/2 the time

QuestionCoveBot:

We will have to multiply: Phenotype \[\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }\] Genotype \[\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }\] So, what is 1/2*1/2?